OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -x + 2y = 8 -x + 10y = 48 I don't just want the answer, I'd really like it if someone walked me through it please

OpenStudy (anonymous):

subtract the equations... that will eliminate the x... what do you have now?

OpenStudy (anonymous):

2y + 10y = 8 + 48 so 12y = 56 ?

OpenStudy (anonymous):

it looks like you add the y's and the # on the right of the equal sign. you need to subtract the equations. subtracting the top minus the bottom equation gives: -8y = -40 can you solve from her?

OpenStudy (anonymous):

***here?

OpenStudy (anonymous):

y = 5?

OpenStudy (anonymous):

once I plugged y in, I got y =5 and x = 2